You've observed the following returns on Crash-n-Burn Computer's stock over the past five years: 7 percent, -14 percent, 20 percent, 32 percent, and 10 percent. Suppose the average inflation rate over this period was 3.61 percent and the average T-bill rate over the period was 4.26 percent. What was the average real risk premium over this time period? [Hint: Step 1: Calculate the average real rate of return for Crash-n-Burn Computer. Step 2: Calculate the average real rate of return for the T-bill. Step 3: Subtract (average real rate of return for Crash-n-Burn Computer) - (average real rate of return for the T-bill). (Enter your answer as a percentage, omit the "%" sign in your response, and round your answer to 2 decimal places. For example, 0.12345 or 12.345% should be entered as 12.35)
Answer #1

Average Nominal Return = [0.07 + (-0.14) + 0.20 + 0.32 + 0.10] / 5
Average Nominal Return = 0.55 / 5
Average Nominal Return = 0.11 or 11.00%

Average Real Return = (Average Nominal Return - Average Inflation Rate) / (1 + Average Inflation Rate)
Average Real Return = (0.1100 - 0.0361) / (1 + 0.0361)
Average Real Return = 0.0739 / 1.0361
Average Real Return = 0.0713 or 7.13%

Average Real Risk-free Rate = (Average Nominal Risk-free Rate - Average Inflation Rate) / (1 + Average Inflation Rate)
Average Real Risk-free Rate = (0.0426 - 0.0361) / (1 + 0.0361)
Average Real Risk-free Rate = 0.0065 / 1.0361
Average Real Risk-free Rate = 0.0063 or 0.63%

Average Real Risk Premium = Average Real Return - Average Real Risk-free Rate
Average Real Risk Premium = 7.13% - 0.63%
Average Real Risk Premium = 6.50%